A company has the following budgeted information:

Product J

Product K

Product L

Sales volume (units)

4,000

3,000

8,000

Unit selling price

$10

$12

$8

Unit variable cost

$8

$9

$7

Budgeted fixed costs for the upcoming period are $150,000.

What is the budgeted break-even in total units?
Answer #1

Total sales = 4000+3000+8000 = 15,000

Sales mix = (4,000/15,000) : (3000/15000) : (8000/15000)

= 27% : 20% : 53%

Weighted average contribution margin

= [(10-8)*27%] + [(12-9)*20%] + [(8-7)*53%)]

= 1.67

Breakeven point = Fixed cost/Weighted-average contribution margin

= 150,000/1.67

= 89,820 units (rounded)